OpenStudy (anonymous):

Plans for a bridge are drawn on a coordinate grid. One girder of the bridge lies on the line y = –9x – 9. A perpendicular brace passes through the point (–1, 7). Write an equation of the line that contains the brace.

OpenStudy (anonymous):

Do they give u an example??

OpenStudy (anonymous):

y – 1 = –1/9(x + 7) x – 7 = –9(y + 1) y – 7 = 1/9( x + 1) y – 7 = –9(x + 1) one of those is the answer

OpenStudy (anonymous):

I think its C...but check with someone else also. I think I wrong